# A Novel Knapsack-based Financial Portfolio Optimization using Quantum Approximate Optimization Algorithm

### v2

@@ Abstract
Keywords: best-known solution (BKS), financial portfolio optimization, knapsack problem, quantum approximate optimization algorithm (QAOA)
@@ 1 Introduction
The financial industry plays a remarkable role in the economic health and growth of a country. In the constantly evolving world of financial markets, technological advancements are shaping traditional portfolio optimization methods, presenting new opportunities and challenges; this process involves enlarging assets to reduce risk by offsetting individual risk profiles crucial to the investment process [1]. Although there are many portfolio optimization models, they possibly have limitations. For example, the mean-variance model, which provides solutions as percentages of the total budget, can result in fractional allocations of nonfeasible assets [2]. In the constant pursuit of optimization, some studies have explored extensions to portfolio optimization [3]. Among these, a technique that reformulates portfolio optimization as a knapsack problem has been proposed [4, 5].
@@
In [2], the knapsack problem was adapted for portfolio optimization by treating all assets included in the portfolio as items. The value or profit of each asset is represented by its expected return, usually estimated using historical data or forecasting techniques. The weight or volume of each item corresponds to a risky asset, typically measured by its standard deviation or the variance-covariance matrix of the assets. The knapsack capacity corresponds to the budget or available capital that can be invested in the portfolio. By reformulating the problem as a knapsack, the primary objective is to maximize the total value or profit of the portfolio while satisfying the capacity constraint and potentially other constraints, such as a target return or a minimum number of assets. However, similar to [6], the knapsack problem is considered NP-hard even with polynomial-bounded weights and values. Therefore, exploring new paradigms for optimization owing to the computational complexity of solving this problem is crucial [7].
@@
Similarly, the paradigm of quantum computing, leveraging the properties of quantum mechanics, has been developed to solve complex problems intractable for classical computers [8]. The financial domain is a primary aspect of quantum computing, with applications spanning price derivation, risk modeling, portfolio optimization, and fraud detection [9]. As technology redefines problem-solving in finance, the quantum approximate optimization algorithm (QAOA), rooted in quantum computing, offers the promise of efficiently finding approximate solutions for computationally demanding problems within the polynomial-bounded NP optimization complexity class. In [10, 11], the application of QAOA to the knapsack problem is explored for optimization purposes.
Herein, we present an approach that harnesses the efficacy of QAOA in addressing the complex problem of portfolio optimization by reformulating it as a knapsack problem. Although additional research is required, our findings offer valuable insights into the following areas.
@@
  • Our proposed model for portfolio selection formulates the problem as a knapsack concern by incorporating the expected return from the Markowitz model and setting the capacity according to the knapsack framework.
@@
  • Our study presents a QAOA algorithm using a quantum walk mixer for the knapsack problem while incorporating a shallow circuit layer to decrease computational complexity and improve solution quality.
  • Our extensive empirical analysis reveals a consistent enhancement in identifying optimal solutions for the knapsack problem, providing an impressive approximation ratio ranging from 100% to 98%, with 2–5 stock selection cases. We also underscore the potential of quantum algorithms as robust, forward-looking solutions for complex financial optimization.
The paper is organized as follows: Section 2 provides an overview of the relevant background work. Section 3 explores the research concerning the proposed method. Section 4 details our proposed method, including the functionality of each step. Section 5 demonstrates the meticulous design and execution of the experiment to obtain the desired results. Finally, Section 6 dissects the findings of the proposed method, underscoring its achievements and limitations while presenting a comprehensive conclusion, key takeaways, and potential future directions.
@@ 2.1 Portfolio Optimization
Portfolio optimization is a mathematical framework that maximizes returns while minimizing risks through strategically selecting assets within an investment portfolio [12]. This is typically achieved by strategically allocating the proportion of each asset in the portfolio to optimize the risk-return tradeoff by considering the specified risk tolerance. The process involves four steps: i) identifying suitable assets, ii) projecting anticipated yields based on historical data for future forecasts, iii) quantifying the risk by assessing the uncertainty of each asset, and iv) selecting the optimal portfolio that maximizes the expected yield for a given risk level [13]. One of the various models used in this study was the Markowitz model, developed by Harry Markowitz in 1952 [14]. This analysis is in conjunction with the variance of the rate of return, providing a significant assessment of portfolio risk under a rational framework of assumptions. The Markowitz model represented the maximum expected yield by allocating funds into stocks as follows [15]:
$R_{i}=\sum_{t=1}^{\infty}d_{it}r_{it}$ (1)
where ${r_{i}}_{t}$ indicates the anticipated return at time $t$ per stock invested in, and ${d_{i}}_{t}$ is the rate at which return in the $i_{th}$ security where time $t$ is discounted back to the presents. The standard deviation—the variance of return—is a statistical measure used as an indicator of the uncertainty or risk linked to return. These statistical indicators effectively measure the extent to which returns deviate unpredictably from the average value over a specific period. The variance represents the degree of variation exhibited by the return $R_{i}$ concerning the expected return $[E(R_{i})]$, as illustrated by the following equation:
$\sigma_{i}^{2}=\frac{1}{N}\sum_{j=1}^{N}[R_{j}-E(R_{j})]^{2}$ (2)
The covariance of returns measures the relative riskiness of a security within a portfolio of securities. For two securities, denoted as $i$ and $j$, the covariance of their returns is defined as
$\sigma_{ij}=E\left\{[R_{i}-E(R_{i})][R_{j}-E(R_{j})]\right\}$ (3)
@@
$\rho_{ij}=\frac{\sigma_{ij}}{\sigma_{i}\sigma_{j}}$ (4)
@@
$\begin{array}[]{l}R_{p}=\sum_{t=1}^{\infty}\sum_{i=0}^{N}d_{it}r_{it}X_{i}\\ \\ =\sum_{i=0}^{N}X_{i}\left(\sum_{t=1}^{\infty}d_{it}r_{it}\right)\\ \\ =\sum_{i=0}^{N}X_{i}R_{i},\end{array}$ (5)
@@
$\sigma_{p}^{2}=\sum_{i=0}^{N}X_{i}^{2}\sigma_{i}^{2}+\sum_{i=0}^{N}\sum_{j=0}^{N}X_{i}X_{j}\sigma_{ij}$ (6)
where $\sigma_{p}^{2}$ is the variance of the portfolio, $X_{i}$ is the percentage of the investor’s assets that are allocated to the $i_{th}$ asset, and the $\sigma_{i}^{2}$ represents the variance of the asset $j$ and the covariance between the returns for assets $i$ and $j$ denoted as $\sigma_{i}{}_{j}$. Reportedly, traditional asset allocation methods, such as the Markowitz theorem, provide solutions in percentages, potentially suggesting the allocation of half of a market share, making it impractical [2]. Therefore, proposing a method for determining the number of shares for each asset is crucial; this involves the conversion of expected returns, prices, and budget into interval values and determines the priority and importance of each share by framing it within a knapsack-based model.
@@ 2.2 Knapsack Problem
Herein, a given set of items with known sizes is selected and packed into a knapsack with a fixed capacity [16]. This problem is one of the simpler NP-hard problems in combinatorial optimization because it focuses on maximizing an objective function while adhering to a single resource constraint. To find the exact solutions, some techniques have been employed, such as relaxations, bounds, reductions, and other algorithmic approaches [17]. These techniques include genetic algorithms [18], dynamic programming [19], simulated annealing [20], Tabu search [21], and greedy algorithm [22]. These classical approaches are instrumental in providing optimal solutions. Although these conventional approaches are successful and valuable, they are limitations in computation in the classical domain. Furthermore, the new paradigm of quantum computing introduces optimization techniques, including QAOA, to tackle problems such as the knapsack problem, yielding better solutions than classical computation techniques [23].
@@ 2.3 Quantum Approximate Optimization Algorithm (QAOA)
In combinatorial optimization, QAOA excels as solutions tailored for quantum computing while leveraging the strengths of classical computing [24, 25]. QAOA, a hybrid quantum-classical algorithm, has demonstrated remarkable effectiveness in addressing recent NP-hard problems, including Max-Cut [26], traveling salesman problem [27], and quadratic unconstrained binary optimization (QUBO) [28]. Given a combinatorial optimization problem involving an N-bit binary string represented as $z=z_{1}\cdot\cdot\cdot z_{N}$, with a classical objective function $f(z):\{0,1\}^{N}\rightarrow\mathbb{R}$ to be maximized [29]. The goal is to find a solution $z$ that satisfies an approximation condition as follows:
$\frac{f(z)}{f_{max}}\geqslant r^{*}$ (7)
where $f_{max}=max_{z}f(z)$, and $r^{*}$ is a desired approximation ratio. Accordingly, the QAOA algorithm tackles this problem by encoding, which involves mapping the classical objective function $f(z)$ into the phase Hamiltonian $H_{c}$ in order to find the optimal eigenvalues:
$H_{c}|z\rangle=f(z)|z\rangle$ (8)
@@
$B=\sum_{j=1}^{N}\sigma_{j}^{x}$ (9)
are applied alternately with controlled durations, generating a wave function:
|$\ket{\psi_{p}(\vec{\gamma},\vec{\beta})}=e^{-i\beta_{p}B}e^{-i\gamma_{p}H_{c}}\cdots e^{-i\beta_{1}B}e^{-i\gamma_{1}H_{c}}\ket{+}^{\otimes N}$ (10)
This variational wave function is parameterized by $2p$ variational parameters, $\gamma$ and $\beta$. The expected value of $H_{c}$ in this variational state is determined through repeated measurements on a computational basis:
$f_{p}(\vec{\gamma},\vec{\beta})=\bra{\psi_{p}(\vec{\gamma},\vec{\beta})}H_{c}\ket{\psi_{p}(\vec{\gamma},\vec{\beta})}$ (11)
@@
$\left(\overrightarrow{\gamma^{*}},\overrightarrow{\beta^{*}}\right)=\arg\max_{\vec{\gamma},\vec{\beta}}f_{p}(\vec{\gamma},\vec{\beta})$ (12)
Next, the approximate ratio showing the QAOA performance is
$r=\frac{f_{p}(\overrightarrow{\gamma^{*}},\overrightarrow{\beta^{*}})}{f_{\text{max}}}$ (13)
Moreover, searching for the approximate ratio is typically performed by starting a random initial estimate of the parameter and performing gradient-based optimization [30].
@@ 3 Related Work
Our study builds upon several key pieces of research that delve into the application of QAOA to the knapsack problem concerning portfolio optimization. QAOA has been employed as an optimization method for portfolio optimization. Herein, the portfolio optimization problem is transformed into a binary version. Accordingly, the weight vector is discretized, with the element taking values of either 0 or 1. In [31], a comprehensive study of quantum computing approaches for multi-napsack problems is proposed by investigating some of the most prominent and state-of-the-art quantum algorithms using different quantum software and hardware tools. Consequently, quantum computing offers the potential for good and fast solutions to multi-knapsack optimization problems in various fields, such as logistics (allocating goods to containers), resource allocation in computing (distributing tasks among different servers), and financial portfolio optimization (allocating assets among different investment opportunities).
In  [32], the Markowitz models of portfolio optimization were converted into binary knapsacks. A hard constraint model was employed by incorporating hard constraints into the quantum algorithm, involving designing mixing operators based on the constraint conditions. Additionally, a combination of $XY$- and $XYY$-mixers was used to encode the constraints in the quantum circuit. $XY$-mixers were used to mix the quantum state and generate a superposition of feasible solutions, whereas $XYY$-mixers were used to enforce the hard constraints. Moreover, using a hard constraint model ensured that the quantum state evolved between feasible solutions satisfying the constraints while allowing for a high degree of flexibility in opting parameters.
Another intriguing study highlights the strengths of the quantum walk optimization algorithm (QWOA) [33] compared to other quantum optimization algorithms while highlighting the challenges posed by the complex and large solution space associated with its lattice structure. Moreover, the quantum mix optimization algorithm (QMOA) was introduced as an extension of QWOA. QMOA enhanced the efficiency of quantum optimization algorithms for portfolio optimization by reducing the number of iterations required for computations [34].

### v3

# Enhancing Knapsack-based Financial Portfolio Optimization Using Quantum Approximate Optimization Algorithm

@@ Abstract
1Department of AI Convergence, Pukyong National University, Nam-gu, Busan 48513, South Korea
2Department of Management Information Systems, Chungbuk National University, 1, Chungdae-ro, Seowon-gu, Cheongju-si, Chungcheongbuk-do, South Korea
*Corresponding author: Youngsun Han (youngsun@pknu.ac.kr)
Keywords
best-known solution (BKS), portfolio optimization, knapsack problem, quantum approximate optimization algorithm (QAOA)
@@ 1 Introduction
The financial industry plays a remarkable role in the economic health and growth of every country. In the constantly evolving world of financial markets, technological advancements are shaping traditional portfolio optimization methods, presenting new opportunities and challenges; this process involves enlarging assets to reduce risk by offsetting individual risk profiles crucial to the investment process [1]. Although there are many portfolio optimization models, they possibly have limitations. For example, the mean-variance model, which provides solutions as percentages of the total budget, can result in fractional allocations of non-feasible assets [2]. In the constant pursuit of optimization, some studies have explored extensions to portfolio optimization [3]. Among these, a technique that reformulates the financial portfolio optimization problem as a knapsack-based problem has been proposed [4, 5].
@@
Knapsack problem [2] was adapted for portfolio optimization by treating all assets included in the portfolio as items. The value of each asset is represented by its expected return, usually estimated using historical data or forecasting techniques. The weight of each item corresponds to a risky asset, typically measured by its standard deviation or the variance-covariance matrix of the assets. The knapsack capacity corresponds to the budget or available capital that can be invested in the portfolio. By reformulating the knapsack problem, the primary objective is to maximize the total value of the portfolio while satisfying the capacity constraint and potentially other constraints, such as a target return or a minimum number of assets. However, similar to [6], the knapsack problem is considered NP-hard even with polynomial-bounded weights and values. Therefore, exploring new paradigms for optimization owing to the computational complexity of solving this problem is crucial [7, 8].
@@
Similarly, the paradigm of quantum computing, leveraging the properties of quantum mechanics, has been developed to solve complex problems intractable for classical computers [9, 10]. The financial domain is a primary aspect of quantum computing, with applications spanning price derivation, risk modeling, portfolio optimization, and fraud detection [11]. As technology redefines problem-solving in finance, the quantum approximate optimization algorithm (QAOA), rooted in quantum computing, offers the promise of efficiently finding approximate solutions for computationally demanding problems within the polynomial-bounded NP optimization complexity class [12]. In [13, 14], the application of QAOA to the knapsack problem is explored for optimization purposes.
In this paper, we reformulate the classical portfolio optimization problem as a knapsack problem that allows for an efficient mapping onto quantum algorithms, particularly advantageous for NP-hard financial optimization. By treating each asset’s expected return as value and its risk as weight, the knapsack model aligns portfolio constraints with total risk, permitting optimization under capacity constraints. Our approach introduces a novel Quantum Walk Mixer (QWM) combined with the Quantum Approximate Optimization Algorithm (QAOA), integrating the Quantum Fourier Transform (QFT) within the oracle. This QWM-QAOA framework ensures feasible states are targeted while reducing computational complexity. Empirical results demonstrate that QWM-QAOA enhances solution space exploration, providing robust portfolio configurations with high approximation ratios even under various noise conditions, thus showcasing a significant advancement in leveraging quantum algorithms for practical finance applications. The contributions of our study can be summarized as follows:
@@
  • We propose a portfolio selection approach that formulates the portfolio problem as a knapsack concern by incorporating expected returns from the Markowitz model [15] and setting the capacity according to the knapsack framework.
@@
  • We present a Quantum Walk Mixer to Quantum Approximate Optimization Algorithm (QWM-QAOA) for the knapsack problem, integrating a shallow circuit layer to decrease computational complexity and improve solution quality.
  • We demonstrate that our model consistently enhances the identification of optimal solutions for the knapsack problem, achieving an impressive approximation ratio ranging from 100% to 95% in scenarios involving the selection of 2 to 8 stocks.
The paper is organized as follows: Section 2 provides the background knowledge. Section 3 explores the related work. Section 4 details our proposed method. Section 5 demonstrates the experimental setup and performance evaluation. Section 6 dissects the findings of the proposed method, underscoring its achievements and limitations. Finally, we conclude our study in Section 7.
@@ 2 Background
In this section, we briefly describe the background knowledge of portfolio optimization, the knapsack problem, and the quantum approximate optimization algorithm (QAOA).
@@ 2.1 Portfolio Optimization
Portfolio optimization is a mathematical framework that maximizes returns while minimizing risks through strategically selecting assets within an investment portfolio [16]. This is typically achieved by strategically allocating the proportion of each asset in the portfolio to optimize the risk-return tradeoff by considering the specified risk tolerance. The process involves four steps: i) identifying suitable assets, ii) projecting anticipated yields based on historical data for future forecasts, iii) quantifying the risk by assessing the uncertainty of each asset, and iv) selecting the optimal portfolio that maximizes the expected yield for a given risk level [17]. One of the various models used in this study was the Markowitz model developed by Harry Markowitz in 1952 [15]. This analysis is in conjunction with the variance of the rate of return, providing a significant assessment of portfolio risk under a rational framework of assumptions. The Markowitz model represented the maximum expected yield by allocating funds into stocks as follows [18]:
$R_{i}=\sum_{t=1}^{\infty}d_{i}{}_{t}r_{i}{}_{t},$ (1)
where ${r_{i}}_{t}$ indicates the anticipated return at time $t$ per stock invested in, and ${d_{i}}_{t}$ is the rate at which return in the $i_{th}$ security where time $t$ is discounted back to the presents. The standard deviation—the variance of return—is a statistical measure used as an indicator of the uncertainty or risk linked to return. These statistical indicators effectively measure the extent to which returns deviate unpredictably from the average value over a specific period. The variance represents the degree of variation exhibited by the return $R_{i}$ concerning the expected return $[E(R_{i})]$, as illustrated:
$\sigma_{i}^{2}=\frac{1}{N}\sum_{i=1}^{N}[R_{i}-E(R_{i})]^{2}.$ (2)
The covariance of returns measures the relative riskiness of a security within a portfolio of securities. For two securities, denoted as $i$ and $j$, the covariance of their returns as
$\sigma_{i}{}_{j}=E\{[R_{i}-E(R_{i})][R_{j}-E(R)]\}.$ (3)
@@
$\rho_{i}{}_{j}=\frac{\sigma_{i}{}_{j}}{\sigma_{i}\sigma_{j}},$ (4)
@@
$\begin{array}[]{l}R_{p}=\sum_{t=1}^{\infty}\sum_{i=0}^{N}d_{i}{}_{t}r_{i}{}_{t}X_{i}\\ \\ =\sum_{i=0}^{N}X_{i}\left(\sum_{t=1}^{\infty}d_{i}{}_{t}r_{i}{}_{t}\right)\\ \\ =\sum_{i=0}^{N}X_{i}R_{i},\end{array}$ (5)
@@
$\sigma_{p}^{2}=\sum_{i=0}^{N}X_{i}^{2}\sigma_{i}^{2}+\sum_{i=0}^{N}\sum_{j=1}^{N}X_{i}X_{j}\sigma_{i}{}_{j},$ (6)
where $\sigma_{p}^{2}$ is the variance of the portfolio, $X_{i}$ is the percentage of the investor’s assets that are allocated to the $i_{th}$ asset, and the $\sigma_{i}^{2}$ represents the variance of the asset $j$ and the covariance between the returns for assets $i$ and $j$ denoted as $\sigma_{i}{}_{j}$. Traditional asset allocation methods, such as the Markowitz theorem, reportedly provide solutions in percentages. This approach can suggest allocating half of a market share, which is often impractical [2]. Therefore, proposing a method for determining the number of shares for each asset is crucial; this involves the conversion of expected returns, prices, and budget into interval values and determines the priority and importance of each share by framing it within a knapsack-based model.
@@ 2.2 Knapsack Problem
In this context, a given set of items with known sizes is selected and packed into a knapsack with a fixed capacity [19]. This problem is one of the simpler NP-hard problems in combinatorial optimization because it focuses on maximizing an objective function while adhering to a single resource constraint. To find the exact solutions, some techniques have been employed, such as relaxations, bounds, reductions, and other algorithmic approaches [20]. These techniques include genetic algorithms [21], dynamic programming [22], simulated annealing [23], Tabu search [24], and greedy algorithm [25]. These classical approaches are instrumental in providing optimal solutions. Although these conventional approaches are successful and valuable, they are limitations in computation in the classical domain. Furthermore, the new paradigm of quantum computing introduces optimization techniques, including QAOA, to tackle problems such as the knapsack problem, yielding better solutions than classical computation techniques [26, 27].
@@ 2.3 Quantum Approximate Optimization Algorithm(QAOA)
In combinatorial optimization, QAOA excels as solutions tailored for quantum computing while leveraging the strengths of classical computing [12, 28]. QAOA, a hybrid quantum-classical algorithm, has demonstrated remarkable effectiveness in addressing recent NP-hard problems, including Max-Cut [29], traveling salesman problem [30], and quadratic unconstrained binary optimization (QUBO) [31]. Consider a combinatorial optimization problem involving an N-bit binary string represented as $z=z_{1}\cdot\cdot\cdot z_{N}$, with a classical objective function $f(z):\{0,1\}^{N}\rightarrow\mathbb{R}$ is maximized. The goal is to find a solution $z$ that provides a high approximation to the maximum values of $f(z)$ [32]. QAOA encodes this classical objective function into a quantum Hamiltonian $H_{c}$,
$H_{c}|z\rangle=f(z)|z\rangle.$ (7)
@@
$B=\sum_{j=1}^{N}{\sigma_{j}^{x}},$ (8)
are applied alternately with controlled duration, resulting in a wave function:
|$\begin{array}[]{l}\ket{\psi_{p}(\vec{\gamma},\vec{\beta})}=e^{-i\beta_{p}B}e^{-i\gamma_{p}Hc}\cdots e^{-i\beta_{1}B}e^{-i\gamma_{1}H_{c}}\ket{+}^{\otimes N}.\end{array}$ (9)
This variational wave function is characterized by $2p$ variational parameters, $\gamma$ and $\beta$. The expected value of $H_{c}$ in this state is determined through repeated measurements on a computational basis:
$f_{p}(\vec{\gamma},\vec{\beta})=\bra{\psi_{p}(\vec{\gamma},\vec{\beta})}H_{c}\ket{\psi_{p}(\vec{\gamma},\vec{\beta})}.$ (10)
@@
$\left(\overrightarrow{\gamma^{*}},\overrightarrow{\beta^{*}}\ \right)\ =\ arg\ \ \underset{\vec{\gamma},\vec{\beta}}{max}\ f_{p}\ (\vec{\gamma},\vec{\beta}).$ (11)
Next, the approximation is assessed by the ratio between the optimized values $f_{p}(\overrightarrow{\gamma^{*}},\overrightarrow{\beta^{*}})$ and the maximum possible value $f_{max}=max_{z}f(z)$, as the approximate ratio:
$r\ =\ \frac{f_{p}\ \left(\overrightarrow{\gamma^{*}},\overrightarrow{\beta^{*}}\ \right)}{f_{max}}.$ (12)
This approximation ratio $r$ reflects how close the QAOA solution is to the optimal classical solution. Typically, the search for this approximate ratio begins with a random initial estimate of the parameter and performing gradient-based optimization [33].
@@ 3 Related Work
Recent studies have explored hybrid approaches to enhance different quantum algorithms for knapsack-based portfolio optimization problems. For example, [34] introduces Quantum-Inspired Evolutionary Algorithm (QIEAs) for difficult knapsack problems. Similarly, [35] used Quantum-inspired meta-heuristic approaches for constrained portfolio optimization problems. These advancements highlight the potential of hybrid quantum-classical approaches for tackling real-world optimization challenges.
Our study builds upon several key pieces of research that delve into the application of QAOA to the knapsack problem concerning portfolio optimization. QAOA has been employed as a portfolio optimization method. In this approach, the portfolio optimization problem is transformed into a binary version. Accordingly, the weight vector is discretized, with the element taking values of either 0 or 1. In [36], a comprehensive study of quantum computing approaches for multi-knapsack problems is proposed by investigating some of the most prominent and state-of-the-art quantum algorithms using different quantum software and hardware tools. Consequently, quantum computing offers the potential for good and fast solutions to multi-knapsack optimization problems in various fields, such as logistics (allocating goods to containers), resource allocation in computing (distributing tasks among different servers), and financial portfolio optimization (allocating assets among different investment opportunities). In [37], the Markowitz models of portfolio optimization were converted into binary knapsacks. A hard constraint model was employed by incorporating hard constraints into the quantum algorithm, involving designing mixing operators based on the constraint conditions. Additionally, a combination of $XY$- and $XYY$-mixers was used to encode the constraints in the quantum circuit. $XY$-mixers were used to mix the quantum state and generate a superposition of feasible solutions, whereas $XYY$-mixers were used to enforce the hard constraints. Moreover, using a hard constraint model ensured that the quantum state evolved between feasible solutions satisfying the constraints while allowing for a high degree of flexibility in opting parameters. Another intriguing study highlights the strengths of the quantum walk optimization algorithm (QWOA) [38] compared to other quantum optimization algorithms while highlighting the challenges posed by the complex and large solution space associated with its lattice structure. Moreover, the quantum mixer optimization algorithm (QMOA) was introduced as an extension of QWOA. It enhanced the efficiency of quantum optimization algorithms for portfolio optimization by reducing the number of iterations required for computations [39].